# Toward efficient resource utilization at edge nodes in federated learning

### v1

@@ Abstract
Federated learning (FL) enables edge nodes to collaboratively contribute to constructing a global model without sharing their data. This is accomplished by devices computing local, private model updates that are then aggregated by a server. However, computational resource constraints and network communication can become a severe bottleneck for larger model sizes typical for deep learning applications. Edge nodes tends to have limited hardware resources (RAM, CPU), and the network bandwidth and reliability at the edge is a concern for scaling federated fleet applications. In this paper, we propose and evaluate a FL strategy inspired by transfer learning in order to reduce resource utilization on devices, as well as the load on the server and network in each global training round. For each local model update, we randomly select layers to train, freezing the remaining part of the model. In doing so, we can reduce both server load and communication costs per round by excluding all untrained layer weights from being transferred to the server. The goal of this study is to empirically explore the potential trade-off between resource utilization on devices and global model convergence under the proposed strategy. We implement the approach using the federated learning framework FEDn. A number of experiments were carried out over different datasets (CIFAR-10, CASA, and IMDB), performing different tasks using different deep-learning model architectures. Our results show that training the model partially can accelerate the training process, efficiently utilizes resources on-device, and reduce the data transmission by around 75% and 53% when we train 25%, and 50% of the model layers, respectively, without harming the resulting global model accuracy. Furthermore, our results demonstrate a negative correlation between the number of participating clients in the training process and the number of layers that need to be trained on each client’s side. As the number of clients increases, there is a decrease in the required number of layers. This observation highlights the potential of the approach, particularly in cross-device use cases.
@@ 1 Introduction
Federated learning (FL) is a privacy-preserving machine learning (ML) training strategy introduced by McMahan et al McMahan et al. (2017). In FL, edge nodes contribute to a global model by locally computing partial model updates, which are then exchanged with a server and combined/aggregated into a global model. By iterating this process, we avoid sharing or transferring private data Konečnỳ et al. (2016a) – instead of moving data to a central server, the model implementation is transferred to the data owners’ local sites, where model training occurs. In this sense, Federated Learning falls in the category of decentralized optimization.
@@
In the most basic FL architecture, a single central server constructs a global model in each communication (training) round by aggregating model parameters sent by the edge nodes. The limited internet connection bandwidth makes the model weight transfer between the edge nodes and server a bottleneck which contributes significantly to the training time of each round Chen et al. (2021). In addition, training a large and complex Neural Network (NN) model at the edge node requires non-trivial time and computational resources (memory, network and CPU). Taken together, it is a challenge to accommodate increasingly complex models in cases where the network connection and the edge hardware have limitations.
@@
Federated learning often involves edge devices that are not homogeneous and have varying computational capabilities. As a result, the research community has focused on addressing the challenges posed by weak computing capacity devices, commonly referred to as stragglers Tandon et al. (2017); Xu et al. (2019). These devices can significantly slow the training process and may need to be excluded from the round. Consequently, training a complex model over massive data in federated settings is challenging due to the presence of stragglers, which can slow down the process and limit the node’s contribution to the round Li et al. (2020).
@@
Therefore, Different approaches have been proposed that can reduce resource utilization during the training of a deep-learning (DL) model, both in centralized settings and decentralized settings. One such strategy is to freeze pre-trained model layers and to add a new output layer to be trained over the new task data, where its weights will be adapted based on the weights of the prior layers (transfer learning or fine-tuning) Too et al. (2019); Liu et al. (2021a). This approach is applicable in a centralized paradigm. For distributed settings, two main parallel strategies are in use to speed up training: 1) replicating a copy of the whole model overall cluster nodes and then using data mini-batches broadcasted by the cluster head (data parallelism) Chen et al. (2018); Zou et al. (2014), and 2) distributing the model’s layers over the cluster nodes, where each layer will be trained using the entire data set, and a cluster coordinator is responsible for parallel communication (model parallelism) Vishnu et al. (2016); Hewett and Grady II (2020); Jia et al. (2018); Shoeybi et al. (2019).
@@
Recently, a new approach inspired by transfer learning aims to selectively freeze a set of layers and only update the remaining layers in each iteration Xiao et al. (2019a). In this paper, we explore the same strategy for local model training in a federated learning setting. The goal is to a) reduce the resource needs on the edge device and, in this way, fit larger models by reducing the training memory and CPU footprint for each update, and b) reduce network transfer costs by reducing the number of updated parameters in each iteration. We are specifically interested in investigating the feasibility of this approach for scenarios where the hardware on the device is too limited to train the entire model effectively (resource starvation). To summarize, the main contributions of this article are:
@@
The rest of this paper is organized as follows. Section 3 provides background on federated learning, model fine-tuning and transfer learning, as well as how the proposed approach relates to general parallel training strategies. Next, we introduce the proposed approach in Section 4. Section 5 describes the experimental settings and the results, followed by a summary of related work in Section 2. Finally, we conclude the paper in Section 6.
2 Related Work
Different approaches have been proposed to distribute the training process workload across a group of machines (Cluster). For large models Dean et. el Dean et al. (2012) have developed a DistBelief framework that supports DL model parallelism. The framework trains large models over a computing cluster with thousands of machines. It comprises two main algorithms, Downpour SGD, responsible for a large number of model replica and adaptive learning rate, while Sandblaster is accountable for the parallelization process. In this study, Jia et al. (2018) authors have investigated the space of parallelization strategies (i.e. SOAP), which includes Sample, Operation, Attribute, and Parameter dimensions to parallelize a DL model. Furthermore, they proposed FlexFlow framework Jia et al. (2018) that uses the SOAP space to search randomly for a fast parallelization strategy for a specific machine. In addition to the frameworks mentioned above, more research has been conducted on model parallelism, such as Sergeev and Del Balso (2018); Park et al. (2020); Gaunt et al. (2017); Chen et al. (2018).
For a large training data sample, Valiant Valiant (1990) has introduced the bulk-synchronous parallel (BSP) model, which parallelises the training process by two main steps. Firstly, place a replica of the whole DL model on each device. Secondly, split the training dataset into mini-batches then distribute them among multiple workers to train the DL model. Finally, each worker synchronises model parameters with a different worker at the end of each iteration Krizhevsky et al. (2012). Moreover, Tensorflow ten (2017), Coffe2 caf (2016), Pytorch pyt (2017) frameworks have been used both data and model parallelisem to parallelise DL training process. Nevertheless, Data parallelism is an efficient technique that can train a small DL model with few parameters. While, in a large model’s case, this becomes an inefficient strategy that caus a scalability jam in large scale distributed training environments.
Chen et al. Chen et al. (2021) were proposed a new scheme named Adaptive Parameter Freezing (APF) tackle the communication bottleneck in federated learning settings. The APF is responsible for freezing and unfreezing converged parameters during the training rounds for intermittent periods. The model was fully trained to identify the stable and unstable parameters for several rounds, and then APF freezes the stable parameters based on threshold increase gradually. In addition, the authors have introduced a mechanism that dynamically adjusts the stability threshold at runtime when most of the parameters have been classified as stable and decreases the stability threshold by one half. The results reveal that this scheme has reduced the communication volume without compromising the model accuracy. However, this approach still relies on memory to cash the prior parameters to check their stability and also requires both CPUs and RAMs to train the entire model at the beginning and each unfreezes period. While our approach train sub-layers of the model selected randomly, which significantly impacts the resources, training time, transferred data, and the final model accuracy, as shown in our results.
Our approach stands out from existing methods in this context due to its unique training strategy in Federated Learning (FL) settings. Our strategy focuses on training different parts of the model every communication round, aiming to achieve several key goals. These goals include, reducing resource utilization, minimizing the communication flow, and enabling the restricted edge devices to participate in the training process without sharing the raw data. Unlike other approaches reported in the literature, where a complete model must be shared to train the model, our approach avoids the need to share a complete model. Instead, we adopt a more efficient technique where each client randomly selects a different model layer every communication round to be trained, which enables stragglers (constrained devices) to effectively participate in each round of the federated training.
3 Federated machine learning and relevant schemes
@@
Federated Machine Learning (FL) is a new solution proposed by google Konečnỳ et al. (2016a, b) to preserve data privacy that aliens with General Data Protection Regulation (GDPR). FL enables users or organizations to collaboratively train a machine learning model without transferring their own data to a central storage system. Instead, the code is moved to the data owners’ local sites in such a paradigm; incremental local updates are combined into a global model every communication round. Typically, a server or aggregator is responsible for managing the client nodes, communication, distributing the global model weight, controlling the training rounds, and generating a global model from received models using FEDAvgMcMahan et al. (2017), FEDProx Sahu et al. (2018) and many other aggregation methods. Firstly, the server starts the training by distributing the ML model architecture with random weight. Secondly, each client will receive the shared model, initiate the training process using local data, and share the weight updates with the server. Third, in each training round, the server combines all obtained weight updates using FEDAvg, or other aggregation methods to update the global model weight. Finally, new weights are distributed across the clients to start a new training round; in this sense, all clients have shared their knowledge without sharing their data Ekmefjord et al. (2021). Hence, FL maintains data privacy, reduces data transferring costs, and shares different client or organization knowledge, especially in healthcare scenarios where the data is more sensitive. In addition, learning from different sources is required to capture more knowledge about a given problem.Yang et al. (2019).
@@ 3.2 Training Parallelization Techniques
Here we discuss the two main approaches that have been introduced to address challenges in large-scale model training. In both cases, the focus is on parallelizing the training process by utilizing large computational resources
@@
    Training deep learning models requires a massive amount of data and computational resources, which can be time-consuming. The data parallelism Valiant (1990) approach has been widely used to address this challenge, using a parameter server architecture to distribute the training workload across multiple workers and speed up the process. In this approach, the training dataset is divided into mini-batches and each worker is assigned a different subset. The parameter server maintains a full copy of the DL model and communicates with workers to synchronize gradients or weights Shallue et al. (2018). Each worker receives a copy of the DL model and a mini-batch of data, computes local gradients, and shares them with the server parameter. The server then updates the model parameters based on the gradients received from the workers, using negative gradient direction or parameter averaging. The latest values are shared with the workers Park et al. (2020); Shazeer et al. (2018). However, the communication channel bandwidth capacity can become a bottleneck as more workers join the training process, leading to slower processing times. To address this issue, a new architecture called AllReduce operations was proposed, which does not depend on the number of workers in the training process to maintain communication channel capacity Thakur et al. (2005).
@@
In this context, we propose a new approach inspired by transfer learning and parallel training to randomly train a deep-learning model based on freezing particular trainable layers randomly. As shown in the next section, this approach reduces the amount of transferred updated local parameters, improving communication efficiency, model training performance, and efficiently utilising the device resources.
@@ 4 Proposed Approach
The approach we propose involves selectively freezing layers during client updates. This technique can be adapted to work with any federated training strategy or aggregation scheme, and it only requires modifications to be made on the client side. As a result, it offers a high degree of flexibility and versatility, making it easy to integrate into existing federated learning workflows. While there are many different aggregation schemes with different properties, we have chosen to exemplify with the Federated Averaging (FedAvg) strategy in this work. FedAvg is an established FL aggregation algorithm proposed in McMahan et al. (2017).
The FedAvg algorithm is a decentralized version of Stochastic Gradient Descent (SGD). In each training round, a subset of $C$ clients receive a copy of the recent global model. These clients execute $E$ epochs over their local dataset $D_{k}$ to update the model weights $W_{k}$. A new global model is constructed once the updated weights $W_{k}$ have been computed and sent back to the server. This new global model is built by averaging all the received updated weights $W_{k}$ contributed by all participating clients $n$ every round.
. This weighted average is then utilized to generate a new version of the global model $M$. This process is repeated until the model converges. The FedAvg algorithm, which is unchanged in this paper, is outlined in Algorithm 1.
@@ 1 Server executes:
2 initialized $W_{0}$ Function FedAVG($k,W_{t-1},W_{t}$):
@@ 4             $S_{t}\leftarrow$ (sample a random set of clients)
6                   $W_{t}^{k}\leftarrow ClientUpdate(k,W_{t},N_{l})$ $W_{t}\leftarrow\sum_{k=1}^{K}\frac{n_{k}}{n}W_{t}^{k}$
@@ 7                   $W_{t}\leftarrow W_{t}(l)-\eta\nabla l(W_{t}(l),b)$
In all experiments that follow we have used the FedAvg implementation in the FEDn federated learning framework Ekmefjord et al. (2021). FEDn is highly scalable and fully distributed and can be used without modification for the evaluated strategy. This also serves to illustrate that the approach can be embedded in a production-grade distributed system in a straightforward manner to reduce resource utilization on the edge device. It should be noted however, that to fully benefit from the potential network transfer reduction, some minor modifications to the FEDn aggregation server would be needed.

### v2

@@ Abstract
Federated learning (FL) enables edge nodes to collaboratively contribute to constructing a global model without sharing their data. This is accomplished by devices computing local, private model updates that are then aggregated by a server. However, computational resource constraints and network communication can become a severe bottleneck for larger model sizes typical for deep learning (DL) applications. Edge nodes tend to have limited hardware resources (RAM, CPU), and the network bandwidth and reliability at the edge is a concern for scaling federated fleet applications. In this paper, we propose and evaluate a FL strategy inspired by transfer learning in order to reduce resource utilization on devices, as well as the load on the server and network in each global training round. For each local model update, we randomly select layers to train, freezing the remaining part of the model. In doing so, we can reduce both server load and communication costs per round by excluding all untrained layer weights from being transferred to the server. The goal of this study is to empirically explore the potential trade-off between resource utilization on devices and global model convergence under the proposed strategy. We implement the approach using the FL framework FEDn. A number of experiments were carried out over different datasets (CIFAR-10, CASA, and IMDB), performing different tasks using different DL model architectures. Our results show that training the model partially can accelerate the training process, efficiently utilizes resources on-device, and reduce the data transmission by around 75% and 53% when we train 25%, and 50% of the model layers, respectively, without harming the resulting global model accuracy. Furthermore, our results demonstrate a negative correlation between the number of participating clients in the training process and the number of layers that need to be trained on each client’s side. As the number of clients increases, there is a decrease in the required number of layers. This observation highlights the potential of the approach, particularly in cross-device use cases.
@@ 1 Introduction
Federated learning (FL) is a privacy-preserving machine learning (ML) training strategy introduced by McMahan et al. McMahan et al. (2017). In FL, edge nodes contribute to a global model by locally computing partial model updates, which are then exchanged with a server and combined/aggregated into a global model. By iterating this process, we avoid sharing or transferring private data Konečnỳ et al. (2016a); Alawadi et al. (2021); Ait-Mlouk et al. (2022) instead of moving data to a central server, the model implementation is transferred to the data owners’ local sites, where model training occurs. In this sense, FL falls in the category of decentralized optimization.
@@
In the most basic FL architecture, a single central server constructs a global model in each communication (training) round by aggregating model parameters sent by the edge nodes. The limited internet connection bandwidth makes the model weight transfer between the edge nodes and server a bottleneck, contributing significantly to each round’s training time Chen et al. (2021). In addition, training a large and complex Neural Network (NN) model at the edge node requires non-trivial time and computational resources (memory, network and CPU). Taken together, it is a challenge to accommodate increasingly complex models in cases where the network connection and the edge hardware have limitations.
@@
FL often involves edge devices that are not homogeneous and have varying computational capabilities. As a result, the research community has focused on addressing the challenges posed by weak computing capacity devices, commonly referred to as stragglers Tandon et al. (2017); Xu et al. (2019). These devices can significantly slow the training process and may need to be excluded from the round. Consequently, training a complex model over massive data in federated settings is challenging due to the presence of stragglers, which can slow down the process and limit the node’s contribution to the round Li et al. (2020).
@@
Therefore, different approaches have been proposed that can reduce resource utilization during the training of a deep learning (DL) model, both in centralized settings and decentralized settings. One such strategy is to freeze pre-trained model layers and to add a new output layer to be trained over the new task data, where its weights will be adapted based on the weights of the prior layers (transfer learning or fine-tuning) Too et al. (2019); Liu et al. (2021a). This approach is applicable in a centralized paradigm. For distributed settings, two main parallel strategies are in use to speed up training: 1) replicating a copy of the whole model overall cluster nodes and then using data mini-batches broadcasted by the cluster head (data parallelism) Chen et al. (2018); Zou et al. (2014), and 2) distributing the model’s layers over the cluster nodes, where each layer will be trained using the entire data set, and a cluster coordinator is responsible for parallel communication (model parallelism) Vishnu et al. (2016); Hewett and Grady II (2020); Jia et al. (2018); Shoeybi et al. (2019).
@@
Recently, a new approach inspired by transfer learning aims to selectively freeze a set of layers and only update the remaining layers in each iteration Xiao et al. (2019a). In this paper, we explore the same strategy for local model training in the FL setting. The goal is to a) reduce the resource needs on the edge device and, in this way, fit larger models by reducing the training memory and CPU footprint for each update, and b) reduce network transfer costs by reducing the number of updated parameters in each iteration. We are specifically interested in investigating the feasibility of this approach for scenarios where the hardware on the device is too limited to train the entire model effectively (resource starvation). To summarize, the main contributions of this article are:
@@
The rest of this paper is organized as follows. Section 2 provides background on FL, model fine-tuning, transfer learning and related work, as well as how the proposed approach relates to general parallel training strategies. Next, we introduce the proposed approach in Section 3. Section 4 describes the experimental settings. Finally, we conclude the paper in Section 5.
2 Background and Related Work
2.1 Federated Learning
@@
FL is a new solution proposed by Google Konečnỳ et al. (2016a, b); Ait-Mlouk et al. (2022) to preserve data privacy that aliens with General Data Protection Regulation (GDPR). FL enables users or organizations to collaboratively train a ML model without transferring their own data to a central storage system. Instead, the code is moved to the data owners’ local sites in such a paradigm; incremental local updates are combined into a global model every communication round Alawadi et al. (2024). Typically, a server or aggregator is responsible for managing the client nodes, communication, distributing the global model weight, controlling the training rounds, and generating a global model from received models using FEDAvgMcMahan et al. (2017), FEDProx Sahu et al. (2018) and many other aggregation methods. Firstly, the server starts the training by distributing the ML model architecture with random weight. Secondly, each client will receive the shared model, initiate the training process using local data, and share the weight updates with the server. Third, in each training round, the server combines all obtained weight updates using FEDAvg, or other aggregation methods to update the global model weight. Finally, new weights are distributed across the clients to start a new training round; in this sense, all clients have shared their knowledge without sharing their data Ekmefjord et al. (2022). Hence, FL maintains data privacy, reduces data transferring costs, and shares different client or organization knowledge, especially in healthcare scenarios where the data is more sensitive. In addition, learning from different sources is required to capture more knowledge about a given problem.Yang et al. (2019).
@@ 2.3 Training Parallelization Techniques
Here, we discuss the two main approaches that have been introduced to address challenges in large-scale model training. In both cases, the focus is on parallelizing the training process by utilizing large computational resources.
@@
    Training DL models requires a massive amount of data and computational resources, which can be time-consuming. The data parallelism Valiant (1990) approach has been widely used to address this challenge, using a parameter server architecture to distribute the training workload across multiple workers and speed up the process. In this approach, the training dataset is divided into mini-batches and each worker is assigned a different subset. The parameter server maintains a full copy of the DL model and communicates with workers to synchronize gradients or weights Shallue et al. (2018). Each worker receives a copy of the DL model and a mini-batch of data, computes local gradients, and shares them with the server parameter. The server then updates the model parameters based on the gradients received from the workers, using negative gradient direction or parameter averaging. The latest values are shared with the workers Park et al. (2020); Shazeer et al. (2018). However, the communication channel bandwidth capacity can become a bottleneck as more workers join the training process, leading to slower processing times. To address this issue, a new architecture called AllReduce operations was proposed, which does not depend on the number of workers in the training process to maintain communication channel capacity Thakur et al. (2005).
@@
2.4 Related Work
Different approaches have been proposed to distribute the training process workload across a group of machines (Cluster). Dean et al. Dean et al. (2012) have developed a DistBelief framework that supports DL model parallelism for large models. The framework trains large models over a computing cluster with thousands of machines. It comprises two main algorithms: Downpour SGD, responsible for a large number of model replicas and adaptive learning rates, while Sandblaster is accountable for the parallelization process. In this study, Jia et al. (2018) authors have investigated the space of parallelization strategies (i.e. SOAP), which includes Sample, Operation, Attribute, and Parameter dimensions to parallelize a DL model. Furthermore, they proposed FlexFlow framework Jia et al. (2018) that uses the SOAP space to search randomly for a fast parallelization strategy for a specific machine. In addition to the frameworks mentioned above, more research has been conducted on model parallelism, such as Sergeev and Del Balso (2018); Park et al. (2020); Gaunt et al. (2017); Chen et al. (2018).
For a large training data sample, Valiant Valiant (1990) has introduced the bulk-synchronous parallel (BSP) model, which parallelises the training process by two main steps. Firstly, a replica of the whole DL model will be placed on each device. Secondly, the training dataset will be split into mini-batches and then distributed among multiple workers to train the DL model. Finally, each worker synchronises model parameters with a different worker at the end of each iteration Krizhevsky et al. (2012). Moreover, Tensorflow ten (2017), Coffe2 caf (2016), and Pytorch pyt (2017) frameworks have been used in both data and model parallelism to parallelise the DL training process. Nevertheless, Data parallelism is an efficient technique that can train a small DL model with few parameters. While in a large model’s case, this becomes an inefficient strategy that causes a scalability jam in large-scale distributed training environments.
Chen et al. Chen et al. (2021) were proposed a new scheme named Adaptive Parameter Freezing (APF) tackle the communication bottleneck in FL settings. The APF is responsible for freezing and unfreezing converged parameters during the training rounds for intermittent periods. The model was fully trained to identify the stable and unstable parameters for several rounds, and then APF freezes the stable parameters based on threshold increase gradually. In addition, the authors have introduced a mechanism that dynamically adjusts the stability threshold at runtime when most of the parameters have been classified as stable and decreases the stability threshold by one-half. The results reveal that this scheme has reduced the communication volume without compromising the model’s accuracy. However, this approach still relies on memory to cash the prior parameters to check their stability and also requires both CPUs and RAMs to train the entire model at the beginning, and each unfreezes period. While our approach train sub-layers of the model selected randomly, which significantly impacts the resources, training time, transferred data, and the final model accuracy, as shown in our results.
Our approach stands out from existing methods in this context due to its unique training strategy in FL settings. Our strategy focuses on training different parts of the model every communication round, aiming to achieve several key goals. These goals include, reducing resource utilization, minimizing the communication flow, and enabling the restricted edge devices to participate in the training process without sharing the raw data. Unlike other approaches reported in the literature, where a complete model must be shared to train the model, our approach avoids the need to share a complete model. Instead, we adopt a more efficient technique where each client randomly selects a different model layer every communication round to be trained, which enables stragglers (constrained devices) to effectively participate in each round of the federated training.
@@ 3 Proposed Approach
The approach we propose involves selectively freezing layers during client updates. Figure 1 shows the abstract diagram of our approach, where each client (i.e. four clients) independently selects a portion of the entire model randomly based on the identified percentage (50% in this diagram, equivalent to five layers out of 10), determines the layers to be trained in FL settings during each communication round. As indicated in the diagram, it’s important to note that each client trains different layers in every round. This technique can be adapted to work with any federated training strategy or aggregation scheme, and it only requires modifications to be made on the client side. As a result, it offers a high degree of flexibility and versatility, making it easy to integrate into existing FL workflows. While there are many different aggregation schemes with different properties, we have chosen to exemplify with the Federated Averaging (FedAvg) strategy in this work. FedAvg is an established FL aggregation algorithm proposed in McMahan et al. (2017).
Figure 1: The abstract diagram depicts the proposed approach for training the ML model with four clients in the FL context. Where each client independently selects 50% of the entire model layers randomly during every training round.
The FedAvg algorithm is a decentralized version of SGD. In each training round, a subset of $C$ clients receive a copy of the recent global model. These clients execute $E$ epochs over their local dataset $D_{k}$ to update the model weights $W_{k}$. A new global model is constructed once the updated weights $W_{k}$ have been computed and sent back to the server. This new global model is built by averaging all the received updated weights $W_{k}$ contributed by all participating clients $n$ every round.
This weighted average is then utilized to generate a new version of the global model $M$. This process is repeated until the model converges. The FedAvg algorithm, which is unchanged in this paper, is outlined in Algorithm 1.
@@ 1 Server executes:
2 initialized $W_{0}$
3 Function FedAVG($k,W_{t-1},W_{t}$):
4
@@ 7            $S_{t}\leftarrow$ (sample a random set of clients)
9                   $W_{t}^{k}\leftarrow ClientUpdate(k,W_{t},N_{l})$
10                   $W_{t}\leftarrow\sum_{k=1}^{K}\frac{n_{k}}{n}W_{t}^{k}$
@@ 9                   $W_{t}\leftarrow W_{t}(l)-\eta\nabla l(W_{t}(l),b)$
In all experiments that follow we have used the FedAvg implementation in the FEDn FL framework Ekmefjord et al. (2022). FEDn is highly scalable and fully distributed and can be used without modification for the evaluated strategy. This also illustrates that the approach can be embedded in a production-grade distributed system in a straightforward manner to reduce resource utilization on the edge device. It should be noted, however, that to fully benefit from the potential network transfer reduction, some minor modifications to the FEDn aggregation server would be needed.